### v1

@@ 2 Technical overview
(see Remark 6). This includes the cubic case, i.e., $k=m=n$, in which its complexity turns out to be similar to that of [20].
@@ 2.1 Preliminaries
Our algorithm will use in a crucial way the notion of dual matrix code and that of Hull. We give both definitions below.
@@
The definition of the hull (Defintion 5) is the only place of the article where this nonstandard bilinear form is used. Besides, every dual or orthogonal complement which appears in the article is taken with respect to the usual inner product
@@
The subsequent lemmas yield two key observations for our algorithms are the following statements which claim that:
@@
$=\operatorname{Tr}(AC)$
$=0.$ (since $C\in h(\mathcal{C})$)
@@ 2.2 Summary of the algorithm
We are given two $k$-dimensional subspaces $\mathcal{C},\mathcal{D}$ of $\mathbb{F}_{q}^{m\times n}$. Our aim is to find two matrices $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ and $Q\in\operatorname{GL}_{n}(\mathbb{F}_{q})$ verifying $\mathcal{D}=P\mathcal{C}Q^{-1}$. If we have found a suitable matrix $P$, computing $Q$ can be done using linear algebra (see Section 5). The strategy for finding $P$ consists first in guessing a pair $(A,B)\in\mathcal{C}^{\perp}\times\mathcal{D}^{\perp}$ such that $B=(P^{-1})^{\top}AQ^{\top}$, that is, a pair $A,B$ which match with respect to the equivalence $\mathcal{D}^{\perp}=(P^{-1})^{\top}C^{\perp}Q^{\top}$ given by Lemma 2. With such a pair at hand one can reduce the equivalence problem to the conjugacy problem of the codes
@@
Indeed, if $B=PAQ^{-1}$ we prove in Lemma 4 further that $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$. Solving a matrix code conjugacy problem in this context is generally as hard as solving $\mathsf{MCE}$ [14], but it is easy in a particular case: when the hull of both codes has dimension 1, the generators of these two hulls are still conjugate. The main two steps in order to find $P$ are the following.
@@
    From $\mathcal{C},\mathcal{D}$, construct two conjugate codes $\mathcal{C}_{A},\mathcal{D}_{B}$ with one-dimensional.
@@
In order to determine these matrices $A$ and $B$, we construct a dictionary whose keys are (normalized and suitably chosen) polynomials $\chi\in\mathbb{F}_{q}[t]$ of degree $m$. The values corresponding to a key $\chi$ are the pairs $(A,U)\in\mathbb{F}_{q}^{m\times n}\times\mathbb{F}_{q}^{m\times m}$ such that the hull $h(\mathcal{C}_{A})$ is one-dimensional and generated by the matrix $U$ with characteristic polynomial $\chi$. We construct this dictionary with roughly $q^{(m-3)/2}$ entries. Then, we apply the same process to $\mathcal{D}$ and look for collisions; the length of the lists ensures that we will find a pair $(A,B)$. This step is explained in detail in Section 3.
@@
Once we have a pair of matrices $(A,B)\in\mathcal{C}^{\perp}\times\mathcal{D}^{\perp}$ such that $h(\mathcal{C}_{A})$ and $h(\mathcal{C}_{B})$ are one-dimensional and generated by conjugate matrices $U$ and $V$, we may easily compute a matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$. We also impose in the collision search that the characteristic polynomial $\chi$ of $U,V$ is squarefree so that $U,V$ are both diagonalizable in an extension of $\mathbb{F}_{q}$. This allows us to reduce to the case of two codes whose one-dimensional hulls are spanned by the same diagonal matrix. In this context, we will observe that the matrix $P$ we are looking for, i.e. the matrix $P$ such that $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$, is the product of $R$ by some invertible matrix which can be expressed as $f(U)$ for some polynomial $f\in\mathbb{F}_{q}[t]$ of degree less than $m$. This polynomial $f$ will be computed by looking at the action of $P$ on a suitable subspace of $\mathcal{C}_{A}$ and $\mathcal{D}_{B}$. This step is explained in Section 4.
@@
This phenomenon could be compared with what happens in classical coding theory, where two problems arise : the permutation equivalence problem (finding a permutation matrix sending a code to another) and the monomial equivalence problem (finding a monomial matrix, i.e. the product of a permutation matrix and a nonsingular diagonal matrix sending one code to another). When the ground field cardinality $q$ is polynomial in the code length, the two problems are known to be polynomially equivalent [27] but Sendrier’s Support Splitting algorithm is on average efficient on the former while being completely inefficient on the latter.
@@ 2.4 Complexity and impact
The complexity of the algorithm is dominated by the collision search in the first step. Given a uniformly random code $\mathcal{C}$ and a uniformly random full-rank matrix $A\in\mathcal{C}^{\perp}$, the codes $\mathcal{C}_{A}$ are uniformly distributed among the matrix codes in $\ker(\operatorname{Tr})\subset\mathbb{F}_{q}^{m\times m}$. Among these, roughly $1/q$ have a one-dimensional hull. By an argument similar to the birthday paradox, computing two lists of length roughly $q^{(m-3)/2}$ is enough to find some collisions. This requires picking matrices $A$, and for each of these, computing the hull of $\mathcal{C}_{A}$. The total time complexity is
$\mathcal{O}\left(\max(k,{k^{\perp}})\ (nm^{\omega-1}+\max(k,{k^{\perp}})m^{2})\ q^{(m-1)/2}\ \max(1,q^{\min(k,{k^{\perp}})-m+1})\right)$
operations in $\mathbb{F}_{q}$, where ${k^{\perp}}\stackrel{{\scriptstyle\text{def}}}{{=}}mn-k$ is the dimension of the dual code $\mathcal{C}^{\perp}$ and $\omega$ is the exponent of matrix multiplication. In the cubic case $m=n=k$, this complexity can be reduced to
$\mathcal{O}\left(n^{4}q^{(n+1)/2}\right).$
@@
$\mathcal{O}\left(m(m+n+1)q^{(m-3)/2}\right)$
elements of $\mathbb{F}_{q}$, which is essentially the size of the computed dictionaries.
@@ 3.1 Structure of the reduction
Lemma 2 shows that the instances $(\mathcal{C},\mathcal{D})$ and $(\mathcal{C}^{\perp},\mathcal{D}^{\perp})$ of $\mathsf{MCE}$ are equivalent. In particular, for complexity reasons, we may switch $(\mathcal{C},\mathcal{D})$ for $(\mathcal{C}^{\perp},\mathcal{D}^{\perp})$: we will systematically choose the instance with the highest dimension. Indeed, since collision search is performed on the dual codes, we fit in the situation where the codes have the smallest possible duals.
Lemma 2 also shows that given $A\in\mathcal{C}^{\perp}$, the matrix $B=(P^{-1})^{\top}AQ^{\top}$ belongs to $\mathcal{D}^{\perp}$. Thee key algorithm lies in the following lemma.
@@
$\mathcal{C}_{A}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{C}A^{\top}\quad\text{and}\quad\mathcal{D}_{B}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{D}B^{\top}.$
@@
This is a straightforward calculation. ∎
@@
By construction the code $\mathcal{C}_{A}\subset\ker(\operatorname{Tr})$. Moreover, in order to have one-dimensional hull, we need the former inclusion to be a strict. Since we require $\mathcal{C}_{A}$ to have the same dimension as $\mathcal{C}$, this means that our algorithm in this form only works for
@@
Allowing ourselves to switch $\mathcal{C}$ and $\mathcal{C}^{\perp}$ at the beginning, it also applies to the case where
${k^{\perp}}<m^{2}-1.$
@@
This yields a list of tuples $(A,B,U,V,R)$ such that the codes $\mathcal{C}_{A}$ and $\mathcal{D}_{B}$ have one-dimensional hulls respectively generated by matrices $U,V$ such that $V=RUR^{-1}$. To these tuples, we then apply the algorithm of Section 4 which allows to find a suitable matrix $P$ such that $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$. In order to compute $Q$, we now need to solve $\mathcal{D}=(P\mathcal{C})\cdot Q$, where $P\mathcal{C}$ is known. This is an easy problem, which is solved by linear algebra as explained in Section 5.
@@
                   $(\chi,U)={\rm ComputeCharpoly}(\mathcal{C}_{A})$
                   if ${\rm gcd}(\chi(t),\chi^{\prime}(t))=1$ then
@@
Data: $k$-dimensional code $\mathcal{D}\subset\mathbb{F}_{q}^{m\times n}$, integer $L$, dictionary $\rm Dict$
Result: List of tuples $(A,B,U,V,R)\in(\mathbb{F}_{q}^{m\times n})^{2}\times(\mathbb{F}_{q}^{m\times m})^{2}\times\operatorname{GL}_{m}(\mathbb{F}_{q})$
where $(A,U)$ is an item in $\operatorname{Dict}$ and $V=RUR^{-1}$ generates $h(\mathcal{D}_{B})$
$\rm CollisionList=[~{}]$
Compute basis of $\mathcal{C}^{\perp}$
while $length(\rm CollisionList)<L$ do
@@
                   $(\chi,V)={\rm ComputeCharpoly}(\mathcal{D}_{B})$
                   if $\chi$ is a key of $\rm Dict$ then
                         for $(A,U)\in\operatorname{Dict}[\chi]$ do
                               Compute a matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$
                               Add $(A,B,U,V,R)$ to $\rm CollisionList$
return $\rm CollisionList$
Algorithm 3 Finding collisions
Algorithm 3: returns list of tuples $(A,U,B,V,R)$ where $h(\mathcal{C}_{A})=\mathbb{F}_{q}U$, $h(\mathcal{C}_{B})=\mathbb{F}_{q}V$ and $V=RUR^{-1}$.
Remark 7
As presented, Algorithm 3 computes a whole list of collisions and then applies to each one the algorithms described in Section 4, until the code equivalence is solved. In practice, one can actually apply the algorithms of Section 4 on the fly each time a collision is found until we find a collision for which the algorithm of Section 4 yield the searched code equivalence.
@@ 3.2 Distribution of the computed matrix spaces and polynomials
Given $\mathcal{D}\in\operatorname{Gr}_{k}(\mathbb{F}_{q}^{m\times m})$ whose elements have trace zero, any element $(\mathcal{C},A)\in\Phi_{|X}^{-1}(\mathcal{D})$ satisfies $A\in\mathcal{C}^{\perp}$. Given a rank $m$ matrix $A\in\mathbb{F}_{q}^{m\times n}$, the codes $\mathcal{C}$ such that $\mathcal{C}A^{\top}=\mathcal{D}$ and $(\mathcal{C},A)\in X$ are exactly the complementary subspaces of $\ker(\phi_{A})$ in $\phi_{A}^{-1}(\mathcal{D})$. The number of elements in $\Phi_{|X}^{-1}(\mathcal{D})$ is the number of rank $m$ matrices in $\mathbb{F}_{q}^{m\times n}$ multiplied by the number of complementary subspaces of an $m(n-m)$-dimensional subspace in an $(k+m(n-m))$-dimensional $\mathbb{F}_{q}$-vector space. The latter number is nonzero and does not depend on a particular choice of $\mathcal{D}$. Hence, the map $f_{1}$ is surjective and equidistributed. ∎
@@
In a code $\mathcal{C}_{A}\subset\ker(\operatorname{Tr})$ any element in the hull of $\mathcal{C}_{A}$ satisfies $\operatorname{Tr}(U^{2})=0$.
@@
This is a direct consequence of the definition of the hull (Definition 5.
@@
Recall that Section 2.1 states that asymptotically, $1/q$ of all matrix codes in $\operatorname{Gr}_{k}(\ker\operatorname{Tr})$ have a one-dimensional hull. Hence, the number of these codes is equivalent to $q^{k(m^{2}-1-k)-1}$. Therefore, for big enough $q$, such codes exist, and the map $f_{2}$ is always surjective.
@@
$U$ $\longmapsto(a_{m-3},\dots,a_{0})\quad\qquad(\star\star\star)$
where $\mathbb{F}_{q}^{\times}$ acts on $\mathbb{F}_{q}^{m-2}$ via $\lambda\diamond(a_{m-3},\dots,a_{0})=(\lambda^{3}a_{m-3},\dots,\lambda^{m}a_{0})$. Its image is $(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-1}\times\{0\}))/\mathbb{F}_{q}^{\times}$. Lemma 8 asserts that any element of the form $(a_{m-3},\dots,u_{0})$ corresponding to a separable polynomial has $\sim q^{m^{2}-m}$ preimages under $f_{3}$.
@@
$\mathrm{Sep}_{q,m}\subset(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-1}\times\{0\}))/\mathbb{F}_{q}^{\times}$
@@
The map $f_{3}$ is asymptotically equidistributed on $f_{3}^{-1}(\rm Sep)$ by Lemma 8. The map $f_{2}$ is equidistributed by Lemma 7, and $f_{1}$ is equidistributed by Lemma 5. ∎

### v2

@@ 2 Technical overview
(see Remark 6). This includes the cubic case, i.e., $k=m=n$, in which the complexity of our algorithm turns out to be similar to that of [20].
@@ 2.1 Preliminaries
Our algorithm will use in a crucial way the notion of dual matrix code and that of Hull. We give both definitions below. Recall that the trace $\operatorname{Tr}(M)$ of a square matrix $M$ is the sum of its diagonal coefficients.
@@
The definition of the hull (Definition 6) is the only place of the article where this nonstandard bilinear form is used. Besides, every dual or orthogonal complement which appears in the article is taken with respect to the usual inner product
@@
The subsequent lemmas yield two key observations for our algorithms:
@@
$=\operatorname{Tr}(AC)=0.$ (since $C\in h(\mathcal{C})$)
@@ 2.2 Summary of the algorithm
We are given two $k$-dimensional subspaces $\mathcal{C},\mathcal{D}$ of $\mathbb{F}_{q}^{m\times n}$. Our aim is to find two matrices $P\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ and $Q\in\operatorname{GL}_{n}(\mathbb{F}_{q})$ verifying $\mathcal{D}=P\mathcal{C}Q^{-1}$. If we have found a suitable matrix $P$, computing $Q$ can be done using linear algebra (see Section 5). The strategy for finding $P$ consists first in guessing a pair $(A,B)\in\mathcal{C}^{\perp}\times\mathcal{D}^{\perp}$ such that $B=(P^{-1})^{\top}AQ^{\top}$, that is, a pair $A,B$ which match with respect to the equivalence $\mathcal{D}^{\perp}=(P^{-1})^{\top}C^{\perp}Q^{\top}$ given by Lemma 2. With such a pair at hand, one can reduce the equivalence problem to the conjugacy problem of the codes
@@
Indeed, if $B=PAQ^{-1}$ we prove in Lemma 4 further that $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$. Solving a matrix code conjugacy problem in this context is generally as hard as solving $\mathsf{MCE}$ [14], but it is easy in a particular case: when the hull of both codes has dimension 1, we may easily find conjugate generators of these two hulls. The two main steps in order to find $P$ are the following.
@@
    From $\mathcal{C},\mathcal{D}$, construct two conjugate codes $\mathcal{C}_{A},\mathcal{D}_{B}$ with one-dimensional hull.
@@
In order to determine these matrices $A$ and $B$, we construct a dictionary whose keys are (normalized and suitably chosen) polynomials $\chi\in\mathbb{F}_{q}[t]$ of degree $m$. The values corresponding to a key $\chi$ are the pairs $(A,U)\in\mathbb{F}_{q}^{m\times n}\times\mathbb{F}_{q}^{m\times m}$ such that the hull $h(\mathcal{C}_{A})$ is one-dimensional and generated by the matrix $U$ with characteristic polynomial $\chi$. Then, we apply the same process to $\mathcal{D}$ and look for collisions. This step is explained in detail in Section 3.
@@
Once we have a pair of matrices $(A,B)\in\mathcal{C}^{\perp}\times\mathcal{D}^{\perp}$ such that $h(\mathcal{C}_{A})$ and $h(\mathcal{C}_{B})$ are one-dimensional and generated by conjugate matrices $U$ and $V$, we may easily compute a matrix $R\in\operatorname{GL}_{m}(\mathbb{F}_{q})$ such that $V=RUR^{-1}$. We also impose in the collision search that the characteristic polynomial $\chi$ of $U,V$ is squarefree so that $U,V$ are both diagonalizable in an extension of $\mathbb{F}_{q}$. In this context, we will observe that the matrix $P$ we are looking for, i.e. the one such that $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$, is the product of $R$ by some invertible matrix which can be expressed as $f(U)$ for some polynomial $f\in\mathbb{F}_{q}[t]$ of degree less than $m$. The calculation of this polynomial is explained in Section 4.
@@
This phenomenon could be compared with what happens in classical coding theory, where two problems arise : the permutation equivalence problem (finding a permutation matrix sending a code to another) and the monomial equivalence problem (finding a monomial matrix, i.e. the product of a permutation matrix and a nonsingular diagonal matrix sending one code to another). When the ground field cardinality $q$ is polynomial in the code length, the two problems are known to be polynomially equivalent [28] but Sendrier’s Support Splitting algorithm [27] is on average efficient on the former while being completely inefficient on the latter.
@@ 2.4 Complexity and impact
The complexity of the algorithm is dominated by the collision search in the first step. Given a uniformly random code $\mathcal{C}$ and a uniformly random full-rank matrix $A\in\mathcal{C}^{\perp}$, the codes $\mathcal{C}_{A}$ are uniformly distributed among the matrix codes in $\ker(\operatorname{Tr})\subset\mathbb{F}_{q}^{m\times m}$. Among these, a proportion of approximately $1/q$ have a one-dimensional hull. By an argument similar to the birthday paradox, computing two lists of length roughly $q^{({k^{\perp}}-2)/2}$ is enough to find some collisions. This requires picking matrices $A$, and for each of these, computing the hull of $\mathcal{C}_{A}$. The total time complexity, given in Theorem 3.1, is
$\widetilde{\mathcal{O}}\left(q^{\max(\frac{{k^{\perp}}}{2},{k^{\perp}}-m+2)}\right)$
operations in $\mathbb{F}_{q}$, where ${k^{\perp}}\stackrel{{\scriptstyle\text{def}}}{{=}}mn-k$ is the dimension of the dual code $\mathcal{C}^{\perp}$. In the cubic case $m=n=k$, by switching $\mathcal{C}$ and $\mathcal{C}^{\perp}$, this complexity can be reduced to
$\widetilde{\mathcal{O}}\left(q^{n/2}\right).$
@@
${\mathcal{O}}\left(({k^{\perp}}+m+1)q^{\min(\frac{{k^{\perp}}}{2}-1,m-3)}\right)$
elements of $\mathbb{F}_{q}$, which is essentially the size of the computed dictionaries. In the cubic case, this amounts to
$\mathcal{O}\left(nq^{\frac{n}{2}-1}\right)$
elements of $\mathbb{F}_{q}$. This reduces the space complexity of [20] by a factor $n^{2}$, which, with the initial cubic parameters of MEDS-128, is about 200.
@@ 3.1 Structure of the reduction
Lemma 2 shows that the instances $(\mathcal{C},\mathcal{D})$ and $(\mathcal{C}^{\perp},\mathcal{D}^{\perp})$ of $\mathsf{MCE}$ are equivalent. In particular, for complexity reasons, we may switch $(\mathcal{C},\mathcal{D})$ for $(\mathcal{C}^{\perp},\mathcal{D}^{\perp})$: we will systematically choose the instance with the highest dimension. Indeed, since collision search is performed on the dual codes, we fit in the situation where the codes have the smallest possible duals. Thus from now on, we suppose
${k^{\perp}}\leqslant k.$
Lemma 2 also shows that given $A\in\mathcal{C}^{\perp}$, the matrix $B=(P^{-1})^{\top}AQ^{\top}$ belongs to $\mathcal{D}^{\perp}$. The key of the algorithm lies in the following lemma.
@@
$\mathcal{C}_{A}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{C}A^{\top}\quad\text{and}\quad\mathcal{D}_{B}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathcal{D}B^{\top}$
@@
This is a straightforward computation. ∎
@@
By construction, there is an inclusion $\mathcal{C}_{A}\subset\ker(\operatorname{Tr})$. Moreover, in order to have a one-dimensional hull, we need this inclusion to be a strict. Since we require $\mathcal{C}_{A}$ to have the same dimension as $\mathcal{C}$, this means that our algorithm in this form only works for
@@
This yields a list of tuples $(A,B,U,V,R)$ such that the codes $\mathcal{C}_{A}$ and $\mathcal{D}_{B}$ have one-dimensional hulls respectively generated by matrices $U,V$ such that $V=RUR^{-1}$. To these tuples, we then apply an algorithm of Section 4 which allows to find a suitable matrix $P$ such that $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$. In order to compute $Q$, we now need to solve $\mathcal{D}=(P\mathcal{C})\cdot Q$, where $P\mathcal{C}$ is known. This is an easy problem, which is solved by linear algebra as explained in Section 5.
@@
                   $(\chi,U)={\rm ComputeNormalizedCharpoly}(\mathcal{C}_{A})$
                   if ${\rm gcd}(\chi(t),\chi^{\prime}(t))=1$ and $\chi\notin\operatorname{Dict}$ then
@@
Data: $k$-dimensional code $\mathcal{D}\subset\mathbb{F}_{q}^{m\times n}$, integer $N$, dictionary $\rm Dict$
Result: Triple $(A,B,P)\in(\mathbb{F}_{q}^{m\times n})^{2}\times(\mathbb{F}_{q}^{m\times m})^{2}\times\operatorname{GL}_{m}(\mathbb{F}_{q})$ s.t. $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$
$i=0$
while $i<N$ do
@@
                   $i=i+1$
                   $(\chi,V)={\rm ComputeNormalizedCharpoly}(\mathcal{D}_{B})$
                   if $\chi$ is a key of $\rm Dict$ with value $A$ then
                         Use one of the Algorithms of Section 4 to deduce $P$
                         if success then
                              return $(A,B,P)$
return $\bot$
Algorithm 3 Finding $P$
Algorithm 3: returns a triple $(A,B,P)$ where $\mathcal{D}_{B}=P\mathcal{C}_{A}P^{-1}$.
@@ 3.2 Distribution of the computed matrix spaces and polynomials
Given $\mathcal{D}\in\operatorname{Gr}_{k}(\mathbb{F}_{q}^{m\times m})$ whose elements have trace zero, any element $(\mathcal{C},A)\in\Phi_{|X}^{-1}(\mathcal{D})$ satisfies $A\in\mathcal{C}^{\perp}$. Given a rank $m$ matrix $A\in\mathbb{F}_{q}^{m\times n}$, the codes $\mathcal{C}$ such that $\mathcal{C}A^{\top}=\mathcal{D}$ and $(\mathcal{C},A)\in X$ are exactly the complementary subspaces of $\ker(\phi_{A})$ in $\phi_{A}^{-1}(\mathcal{D})$. The number of elements in $\Phi_{|X}^{-1}(\mathcal{D})$ is the number of rank $m$ matrices in $\mathbb{F}_{q}^{m\times n}$ multiplied by the number of complementary subspaces of an $m(n-m)$-dimensional subspace in a $(k+m(n-m))$-dimensional $\mathbb{F}_{q}$-vector space. The latter number is nonzero and does not depend on a particular choice of $\mathcal{D}$. Hence, the map $f_{1}$ is surjective and equidistributed. ∎
@@
In a code $\mathcal{C}_{A}\subset\ker(\operatorname{Tr})$, any element in the hull of $\mathcal{C}_{A}$ satisfies $\operatorname{Tr}(U^{2})=0$.
@@
This is a direct consequence of the definition of the hull (Definition 6).
@@
Recall that Section 2.1 states that asymptotically, $1/q$ of all matrix codes in $\operatorname{Gr}_{k}(\ker\operatorname{Tr})$ have a one-dimensional hull. Hence, the number of these codes is equivalent to $q^{k(m^{2}-1-k)-1}$ when $q\rightarrow\infty$. Therefore, for big enough $q$, such codes exist, and the map $f_{2}$ is always surjective.
@@
[…]
where $\mathbb{F}_{q}^{\times}$ acts on $\mathbb{F}_{q}^{m-2}$ via $\lambda\diamond(a_{m-3},\dots,a_{0})=(\lambda^{3}a_{m-3},\dots,\lambda^{m}a_{0})$. Its image is $(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-3}\times\{0\}))/\mathbb{F}_{q}^{\times}$. Lemma 8 asserts that any element of the form $(a_{m-3},\dots,u_{0})$ corresponding to a separable polynomial has $\sim q^{m^{2}-m}$ preimages under $f_{3}$.
@@
$\mathrm{Sep}_{q,m}\subset(\mathbb{F}_{q}^{m-2}-(\mathbb{F}_{q}^{m-3}\times\{0\}))/\mathbb{F}_{q}^{\times}$
@@
It is a direct consequence of Lemmas 5, 7 and 8. ∎